find the co-ordinates of the point on y-axis which is at a distance of 5root2 from the point P(3,-2,5)

Grade:

1 Answers

Vasantha Kumari
9 years ago
From the given data, we are clear that the point is on y-axis and its value is 0 for all x.
N ow, the distance between the two points is similar to pythagoras theorem and we can find the third side as given below.
(x2 – x1)2 + (y2 – y1)2=D2
(3–0)2+ (– 2.5–y1)2 = (5v2)2
9 + 6.25+5y1+y12=100
y12 +5y1– 84.75 =0
Solving the given equation gives the co-ordinates of the required point.

Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty